OpenStudy (anonymous):

A student estimated the volume of a liquid in a beaker to be 450 mL. When she poured the liquid into a graduated cylinder, she noted the measurement to be 390 mL. Considering the measurement in the graduated cylinder to be the accepted value, what is the percent error of the estimated volume from the beaker? A. 0.15% error B. 1.5% error C. 15% error D. 154% error

OpenStudy (anonymous):

1% error would be a difference of 4.5mL. Your difference is 60mL, so therefore, the error should be 60/4.5= 13.333% which isn't one of the choices..... Or you could do (60/450) = .13333333 So maybe I'm wrong or 15% is the correct one.

OpenStudy (ghazi):

1-(390/450)*100 approximately = 15 %

OpenStudy (anonymous):

Still 13.333%....